# Prox-DBRO-VR: A Unified Analysis on Decentralized Byzantine-Resilient Composite Stochastic Optimization with Variance Reduction and Non-Asymptotic Convergence Rates

### v6

@@ II-B Problem Statement
A network of $m$ agents connect with each other over an undirected network $\mathcal{G}=\left({\mathcal{R}\cup\mathcal{B},\mathcal{E}}\right)$, where $\mathcal{R}$ ($0\leq\left|\mathcal{R}\right|\leq m$) and $\mathcal{B}$ indicate the sets of reliable and Byzantine agents, respectively, and $\mathcal{E}$ represents the set of undirected communication edges among all agents. The mutual target of all reliable agents is to minimize (min) a general decentralized composite optimization problem as follows:
$\mathop{\min}\limits_{\tilde{x}\in{\mathbb{R}^{n}}}\sum\limits_{i\in\mathcal{R}}{{f_{i}}\left({\tilde{x}}\right)+{g}\left({\tilde{x}}\right)},$ (1)
where $\tilde{x}$ is the decision variable, and ${f_{i}}:{\mathbb{R}^{n}}\to\mathbb{R}$ and ${g}:{\mathbb{R}^{n}}\to\mathbb{R}$, $i\in\mathcal{R}$, are two different objective functions. The local objective function $f_{i}$ can be further decomposed as ${f_{i}}\left({\tilde{x}}\right)=\sum\nolimits_{l=1}^{{q_{i}}}{f_{i}^{l}\left({\tilde{x}}\right)}/q_{i}$, while $g$ features a shared non-smooth part similar to literature [36, 35, 38]. We assume that the optimal solution to (1) exists, denoted by $\tilde{x}^{*}$, and the local sample set associated with agent $i$ as ${\mathcal{Q}_{i}}=\left\{{1,2,\ldots,{q_{i}}}\right\}$, $\forall i\in\mathcal{R}$. To specify the problem, we need to make the following assumptions.
@@
a) For $i\in\mathcal{R}$, the local objective function $f_{i}$ is $\mu_{i}$-strongly convex, and the local component objective function $f_{i}^{l}$, is $L_{i}$-smooth, $\forall l\in{\mathcal{Q}_{i}}$, i.e., $\forall{\tilde{x}},{\tilde{z}}\in{\mathbb{R}^{n}}$,
@@
${\left\|{\nabla f_{i}^{l}\left({\tilde{x}}\right)-\nabla f_{i}^{l}\left({\tilde{z}}\right)}\right\|_{2}}\leq L_{i}{\left\|{\tilde{x}-\tilde{z}}\right\|_{2}},$ (2b)
where $\mu:=\mathop{\min}\nolimits_{i\in\mathcal{R}}\left\{{{\mu_{i}}}\right\}$ and $L:={\max_{i\in\mathcal{R}}}\left\{{{L_{i}}}\right\}$, with ${\kappa_{f}}:=L/\mu$;
@@
Assumption 1-a) is standard in recent literature [16, 15, 13, 12, 14, 10]. According to [44, Chapter 3], we know that $0<\mu\leq L$, which indicates $\kappa_{f}\geq 1$. Moreover, in view of (2b), it is not difficult to verify that the local objective function $f_{i}$, $i\in\mathcal{R}$, is $L$-smooth as well. Under Assumption 1, the optimal solution $\tilde{x}^{*}$ to (1) exists uniquely. The consideration of the possibly non-smooth term $g$ is meaningful, which finds substantial applications in various fields, such as the standard 1-norm regularization term in sparse machine learning [35, 36], the non-smooth indicator function in model predictive control [5], and the non-differentiable emission cost in energy resource coordination over smart grids [12].
The following assumption is also standard in recent literature [28, 27, 33].
@@
(Network Connectivity). All reliable agents form a static network, denoted as $\left({\mathcal{R},{\mathcal{E}_{\mathcal{R}}}}\right)$, which is bidirectionally connected.
@@ II-C Problem Reformulation
To guarantee all reliable agents reach a consensus at the optimal solution, we need to reformulate (1) into an equivalent consensus problem. To achieve this goal, a global decision vector $x=\left[{x_{1}^{\top},x_{2}^{\top},\ldots,x_{\left|\mathcal{R}\right|}^{\top}}\right]^{\top}\in{\mathbb{R}^{\left|\mathcal{R}\right|n}}$ with $\left|\mathcal{R}\right|$ local copies of the decision variable $\tilde{x}$, is introduced, subject to (s.t.) the consensus constraint ${x_{i}}={x_{j}},\left({i,j}\right)\in\mathcal{E}$. Therefore, it is natural to rewrite (1) as
$\mathop{\min}\limits_{x\in{\mathbb{R}^{\left|\mathcal{R}\right|n}}}$ $F\left(x\right)+G\left(x\right),\hfill$ (3)
@@
II-D Robust Consensus Problem Setup
To enhance the robustness of the consensual aggregation process, we extend a scalar-valued consensus technique designed in [45] to its vector-valued domain. We can solve for the globally optimal solution ${x^{*}}$ of (3) via the following norm-penalized approximation
${x^{*}}:=\arg\mathop{\min}\limits_{x}\sum\limits_{i\in\mathcal{R}}{\left({{f_{i}}\left({{x_{i}}}\right)+g\left({{x_{i}}}\right)+{\phi_{i}}\sum\limits_{j\in{\mathcal{R}_{i}}}{{{\left\|{{x_{i}}-{x_{j}}}\right\|}_{a}}}}\right)},$ (4)
where $a\geq 1$, and $\phi_{i}$ is the local uncoordinated penalty parameter associated with each reliable agent $i$, $\forall i\in\mathcal{R}$. The introduction of the total-variation norm penalty provides a resilient replacement of the consensus constraint, i.e., the controllable distance between $x_{i}$ and $x_{j}$. The distance is controlled by the uncoordinated penalty parameter $\phi_{i}$, which means that a larger $\phi_{i}$ can bring a small gap between $x_{i}$ and $x_{j}$, $\left({i,j}\right)\in{\mathcal{E}_{\mathcal{R}}}$. To a certain extent, (1) can be considered as a soft relaxation of (4), because the former tolerates the dissimilarity among agents, for instance, the disagreement between reliable agents and Byzantine agents. The equivalence between the norm-penalized approximation problem (4) and the original optimization problem (1) is proved in Theorem 1.
@@ III Algorithm Development
III-A Connection with Existing Works
Lian et al. in [46] design a decentralized stochastic gradient descent algorithm, namely DPSGD, to resolve efficiently the transformed optimization problem (3), in an ideal situation. The ideal situation fails to consider the presence of any malfunctioning or malicious agents, which may not be avoided in practical applications [24, 23, 47, 26, 22, 20, 9, 21]. We next find out the reason why DPSGD cannot be applied directly to solving (3) when there are Byzantine agents in the network, and then seek out a feasible improvement, based on DPSGD, to maintain Byzantine resilience. We first recap the updates of the generalized DPSGD as follows:
@@ III-A Connection with Existing Works
${x_{i,k+1}}=\sum\limits_{j\in{\mathcal{N}_{i}}\cup\left\{i\right\}}{{w_{ij}}{{\bar{x}}_{j,k}}},$ (5b)
where ${{\cal N}_{i}}={{\cal R}_{i}}\cup{{\cal B}_{i}}$, $\alpha_{k}$ denotes a constant or decaying step-size, ${\nabla f\left({{x_{i,k}}}\right)}$ is the local batch gradient, ${w_{ij}}$ is the $i$-th row and $j$-th column element of a doubly stochastic weight matrix, meeting $\sum\nolimits_{j\in{{\cal N}_{i}}}{{w_{ij}}}=\sum\nolimits_{j\in{{\cal N}_{i}}}{{w_{ji}}}=1$. If there is a Byzantine agent $o$ with one reliable neighbor $i$, ${{\tilde{x}}_{io,k}}$ could be an incorrect or misleading information (depending on whether agent $o$ is out of action or manipulated by adversaries), to its reliable neighboring agents, at $k$-th iteration. Then, ${x_{i,k+1}}$ can be arbitrarily deviate from its true model, if Byzantine agent $o$ is manipulated, since the adversary may send a maliciously falsifying message to agent $i$. For instance, Byzantine agent $o$, $o\in{{\cal B}_{i}}$, can blow $x_{i,k+1}$ up to infinity through transmitting continually a vector with infinite elements to its reliable neighbor $i$. Another example is that Byzantine agent $o$ can deter all reliable agents from achieving consensus at iteration $k$, via sending various values $\tilde{x}_{io,k}$ to its different reliable neighboring agents $i\in{\mathcal{R}_{o}}$. The main reason for the above mentioned issues comes from the fact that the aggregation step (5b) is rather vulnerable to Byzantine problems. In fact, similar problems also prevail in decentralized work [12, 10, 43, 1, 16, 13, 14, 15, 11, 2]. Therefore, the SGD family contains two important extensions, RSA [28] and [27], both of which achieve Byzantine resilience based on a robust consensus method [45]. [27] is a decentralized extension of RSA [28]. The theoretical analysis of both RSA [28] and [27] is based on a bounded-variance assumption on the local stochastic gradient. With this assumption and the other standard assumptions (see [27] for details), the sequence ${\left\{{{x_{k}}}\right\}_{k\geq 0}}$ generated by the decentralized algorithm proposed in [27] takes a convergent form of
@@
where $\eta$ is a positive constant satisfying $10<\eta{\alpha_{k}}<$ 1, ${\Delta_{0}}:=\sum\nolimits_{i\in\mathcal{R}}{32n{\phi^{2}}{{\left|{{\mathcal{R}_{i}}}\right|}^{2}}}+4n{\phi^{2}}{\left|{{\mathcal{B}_{i}}}\right|^{2}}+2\sigma_{i}^{2}$ (${\sigma_{i}}>0$ is the bounded variance yielded by the biased evaluation of the local batch gradients) and ${\Delta_{1}}:=\sum\nolimits_{i\in\mathcal{R}}{n{\phi^{2}}{{\left|{{\mathcal{B}_{i}}}\right|}^{2}}/\gamma}$. Based on (6), one can establish either a sub-linear convergence rate of a convergent error determined by the number of Byzantine agents, or a linear convergence rate with a convergent error determined jointly by the number of Byzantine agents and reliable agents, together with the bounded variance ${\sigma_{i}}$. In fact, this bounded variance ${\sigma_{i}}$ exists commonly in recent literature, such as [43, 28, 27, 14]. Therefore, this paper aims to reduce asymptotically this bounded variance in the linear convergence result and discard the bounded-variance assumption as well. Inspired by the recent exploration of decentralized variance-reduced stochastic gradient algorithms diffusion-AVRG [2], S-DIGing [15], GT-SAGA/GT-SVRG [11], GT-SARAH [1] and Push-LSVRG-UP [16] that seek the solution to the optimization problem under an ideal Byzantine-free situation, we consider introducing two popular localized variance-reduction techniques SAGA [41] and LSVRG [42] to reduce asymptotically the variance arising in the course of evaluating the noisy stochastic gradients.
@@ III-B A General Algorithmic Framework
Based on the above analysis, we propose a decentralized Byzantine-resilient stochastic-gradient algorithmic framework in Algorithm 1 to resolve (4) in the presence of Byzantine agents.
Algorithm 1 Prox-DBRO-VR Framework
0:  Each reliable agent $i$, $i\in\mathcal{R}$, initializes with an arbitrary starting point $x_{i,0}\in{\mathbb{R}^{n}}$, a well-designed constant or decaying step-size $\alpha_{k}>0$, and the proper uncoordinated penalty parameters ${{\phi_{i}}}>0$, $\forall i\in\mathcal{R}$.
1:  for all $k=0,1,2,\ldots$ do
2:     Each reliable agent $i$, $i\in\mathcal{R}$, sends its current local model ${x_{i,k}}$ to its neighbors $j\in{\mathcal{N}_{i}}$ and receives the reliable information ${x_{j,k}}$ or malicious information ${z_{ij,k}}$ from its neighbors $j\in{\mathcal{N}_{i}}$.
3:     Each reliable agent $i$, $i\in\mathcal{R}$, evaluates the local stochastic gradient $r_{i,k}$.
4:     Each reliable agent $i$, $i\in\mathcal{R}$, updates an intermediate variable according to the robust local stochastic gradient descent step:
${\bar{x}_{i,k}}={x_{i,k}}-{\alpha_{k}}r_{i,k}-{\alpha_{k}}{\phi_{i}}\sum\limits_{j\in{\mathcal{N}_{i}}}{{\partial_{{x_{i}}}}{{\left\|{{x_{i,k}}-v_{ij,k}}\right\|}_{a}}},$
[…]
5:     Each reliable agent $i$, $i\in\mathcal{R}$, updates its current local model according to the local proximal mapping step:
${x_{i,k+1}}=\mathop{\arg\min}\limits_{\tilde{x}\in{\mathbb{R}^{n}}}\left\{{g\left({\tilde{x}}\right)+\frac{1}{{2{\alpha_{k}}}}{{\left\|{\tilde{x}-{{\bar{x}}_{i,k}}}\right\|}_{2}^{2}}}\right\}.$
6:  end for

### v12

@@ II-B Problem Statement
A network of $m$ agents connect with each other over an undirected network $\mathcal{G}:=\left({\mathcal{R}\cup\mathcal{B},\mathcal{E}}\right)$, where $\mathcal{R}$ ($2\leq\left|\mathcal{R}\right|\leq m$) and $\mathcal{B}$ indicate the sets of reliable and Byzantine agents, respectively, and $\mathcal{E}$ represents the set of undirected (communication) edges among all agents. We represent the information flow from agent $j$ to its neighboring agent $i$ by the edge $(i,j)\in\mathcal{E}$, with the convention that $i<j$ always holds for $i\in\mathcal{V},j\in{\mathcal{N}_{i}}$. The mutual target of all reliable agents is to minimize (min) a general decentralized composite ERM problem as follows:
$\mathop{\min}\limits_{\tilde{x}}\sum\limits_{i\in\mathcal{R}}{{f_{i}}\left({\tilde{x}}\right)+{g}\left({\tilde{x}}\right)},$ (1)
where $\tilde{x}\in{\mathbb{R}^{n}}$ is the decision variable, and ${f_{i}}:{\mathbb{R}^{n}}\to\mathbb{R}$ and ${g}:{\mathbb{R}^{n}}\to\mathbb{R}$, $i\in\mathcal{R}$, are two different objective functions. The local objective function $f_{i}$ can be further decomposed as ${f_{i}}\left({\tilde{x}}\right)=\sum\nolimits_{l=1}^{{q_{i}}}{f_{i}^{l}\left({\tilde{x}}\right)}/q_{i}$, while the function $g$ serves as a shared nonsmooth objective among all reliable agents similar to literature [38, 37, 41]. Based on the above definitions, problem (1) is also recognized as a composite ERM problem in the presence of Byzantine agents. We assume that the optimal solution to (1) exists, denoted by $\tilde{x}^{*}$, and the local sample set associated with agent $i$ as ${\mathcal{Q}_{i}}=\left\{{1,2,\ldots,{q_{i}}}\right\}$, $\forall i\in\mathcal{R}$. This paper aims to resolve the general composite ERM problem (1) under Byzantine attacks or failures, including but not limited to the zero-sum attack [31], Gaussian attack [26], and same-value attack [27], which will be testified in numerical experiments. In fact, failures, such as agent breakdown and possible disconnection of communication links, can be deemed as a category of not malicious Byzantine problems. We next specify the studied problem via the following standard assumptions.
@@
a) For $i\in\mathcal{R}$, the local objective function $f_{i}$ is assumed to be $\mu_{i}$-strongly convex, and the local component objective function $f_{i}^{l}$ is $L_{i}^{l}$-smooth, $\forall l\in{\mathcal{Q}_{i}}$, i.e., $\forall{\tilde{x}},{\tilde{z}}\in{\mathbb{R}^{n}}$,
@@
${\left\|{\nabla f_{i}^{l}\left({\tilde{x}}\right)-\nabla f_{i}^{l}\left({\tilde{z}}\right)}\right\|_{2}}\leq L_{i}^{l}{\left\|{\tilde{x}-\tilde{z}}\right\|_{2}},$ (2b)
where ${\mu_{i}}$ and $L_{i}^{l}$ are the strongly-convex and smooth parameters, respectively.
@@
For convenience, we define $\mu:=\mathop{\min}\nolimits_{i\in\mathcal{R}}\left\{{{\mu_{i}}}\right\}$ and $L:={\max_{i\in\mathcal{R}}}\left\{{{L_{i}}}\right\}$ with ${L_{i}}:=\sum\nolimits_{l=1}^{{q_{i}}}{L_{i}^{l}}/{q_{i}}$. It can be verified that the global objective function is $L$-smooth and $\mu$-strongly convex, with a condition number denoted by ${\kappa_{f}}:=L/\mu$.
@@
Assumption 1-a) is standard in recent literature [15, 13, 12, 14, 10]. According to [45, Chapter 3], we know that $0<\mu\leq L$, which indicates $\kappa_{f}\geq 1$. Moreover, in view of (2b), it can be verified that the local objective functions $f_{i}$, $i\in\mathcal{R}$, are $L$-smooth as well. Under Assumption 1, the optimal solution $\tilde{x}^{*}$ to (1) exists uniquely. The consideration of the possibly nonsmooth term $g$ is meaningful, which finds substantial applications in various fields, such as the standard 1-norm regularization term in sparse machine learning [39, 37, 38], a nonsmooth indicator function in model predictive control [5] to handle equality and set constraints, and a nonsmooth indicator function in energy resource coordination [12] to handle inequality and set constraints.
@@
(Network connectivity) All reliable agents form a static network, denoted as ${{\cal G}_{\cal R}}:=\left({\mathcal{R},{\mathcal{E}_{\mathcal{R}}}}\right)$, which is bidirectionally connected.
Remark 2
Assumption 2 is standard in recent literature [46, 34, 31, 20], which implies that each reliable agent must have at least one reliable neighbor and an arbitrary number of Byzantine neighbors to enable the communication with any other reliable agents in the network. There are many examples satisfying Assumption 2 and a straightforward instance is the full-connected network among all agents. From another perspective, there must be two reliable agents that fail to exchange messages if ${{\cal G}_{\cal R}}$ is disconnected. In this scenario, the disconnected agent is also judged as a Byzantine agent subject to possible failures. Assumption 2 imposes few restrictions on the reliable network ${{\cal G}_{\cal R}}$ than the pioneering literature [17], which requires at least $\left|\mathcal{B}\right|+1$ paths between any two reliable agents when there are $\left|\mathcal{B}\right|$ Byzantine agents in the network. In view of this, there are some network examples, for instance the Dumbbell network [32] satisfying Assumption 2 but violating the resilient network assumption made in [17].
@@ II-C Problem Reformulation
To guarantee all reliable agents reach a consensus at the optimal solution, we need to reformulate (1) into an equivalent consensus problem. To achieve this goal, a global decision vector $x={\mathbf{col}}{\left\{{{x_{i}}}\right\}_{i\in\mathcal{R}}}\in{\mathbb{R}^{\left|\mathcal{R}\right|n}}$ containing $\left|\mathcal{R}\right|$ local copies of the decision variable $\tilde{x}$, is introduced, subject to (s.t.) the consensus constraint ${x_{i}}={x_{j}},\left({i,j}\right)\in\mathcal{E}_{\mathcal{R}}$. Therefore, it is natural to rewrite (2) as
$\mathop{\min}\limits_{x}F\left(x\right)+G\left(x\right),\hfill$ (3)
@@
II-D Resilient Consensus Problem Setup
To facilitate the design of the resilient aggregation mechanism, we examine a norm-penalized approximation variant of the consensus problem (3), originally studied in [36], as follows:
${x^{*}}:=\arg\mathop{\min}\limits_{x}\sum\limits_{i\in\mathcal{R}}{({{f_{i}}\left({{x_{i}}}\right)+g\left({{x_{i}}}\right)+\frac{\phi}{2}\sum\limits_{j\in{\mathcal{R}_{i}}}{{{\left\|{{x_{i}}-{x_{j}}}\right\|}_{a}}}})},$ (4)
where $a\geq 1$, $\phi$ is the penalty parameter, and ${\mathcal{R}_{i}}$ denotes the set of reliable neighbors of agent $i$. The norm penalty takes a soft-approximation replacement of the consensus constraint, thereby admitting a resilient distance between any two neighboring states. The distance is controlled by the penalty parameter $\phi$, indicating that a larger $\phi$ can bring a smaller gap between $x_{i}$ and $x_{j}$, $\left({i,j}\right)\in{\mathcal{E}_{\mathcal{R}}}$. Therefore, the problem formulation (4) is regarded as a soft approximation of (3) as the former tolerates the state dissimilarities among any neighboring agents, which exhibits resilience to a prevalent issue of data silos in decentralized learning tasks. The equivalence between the soft-approximation problem (4) and the consensus problem (3) with respect to the original problem (1) is proved in Theorem 1.
@@ III Algorithm Development
III-A Connection to Existing Work
Lian et al. in [47] design a decentralized stochastic gradient descent algorithm, namely D-PSGD, to resolve efficiently the consensus problem (3) in an ideal situation. The ideal situation fails to consider the presence of any malfunctioning or malicious agents, i.e., Byzantine agents, which may not be avoided in practical applications [24, 23, 48, 26, 22, 9, 21]. We next find out the reason why D-PSGD cannot be applied directly to solving (3) when there are Byzantine agents in the network, and then seek out a feasible improvement to secure Byzantine resilience. We first recap the updates of the generalized D-PSGD as follows:
@@ III-A Connection to Existing Work
${x_{i,k+1}}=\sum\limits_{j\in{{{\cal R}_{i}}\cup{{\mathcal{B}}_{i}}}}{{w_{ij}}{{{v_{ij,k}}}}},$ (5b)
where ${\mathcal{B}_{i}}$ denotes the set of Byzantine neighbors of agent $i$, $i\in\mathcal{R}$. ${v_{ij,k}}:=\left\{\begin{array}[]{l}\!\!\!{{\bar{x}}_{j,k}},j\in{{\mathcal{R}}_{i}}\\ \!\!\!{z_{ij,k}},j\in{{\mathcal{B}}_{i}}\end{array}\right.$ with ${z_{ij,k}}$ denoted as untrue or misleading information sent by Byzantine agent $j$, $j\in{\mathcal{B}}_{i}$, $\alpha_{k}$ denotes a proper constant or decaying step-size, ${\nabla f\left({{x_{i,k}}}\right)}$ is the local batch gradient, ${w_{ij}}$ is the $i$-th row and $j$-th column element of a doubly stochastic weight matrix s.t. $\sum\nolimits_{j\in{{\mathcal{N}}_{i}}}{{w_{ij}}}=\sum\nolimits_{j\in{{\mathcal{N}}_{i}}}{{w_{ji}}}=1$. Note that both $\mathcal{R}_{i}$ and $\mathcal{B}_{i}$ exclude agent $i$ (itself). If there is a reliable agent $i$ with a Byzantine neighbor $b$, $b\in{{\mathcal{B}}_{i}}$, then ${{z}_{ib,k}}$ could be untrue or misleading information (depending on whether agent $b$ experiences malfunction or manipulated by adversaries), to its reliable neighboring agents at $k$-th iteration. Since any Byzantine agent is assumed to be omniscient and able to learn from update rules, if agent $b$ sends an elaborately falsified message to its reliable neighbor $i$, then ${x_{i,k+1}}$ could arbitrarily deviate from its true model. For instance, Byzantine agent $b$ can blow $x_{i,k+1}$ up to infinity through continually transmitting a vector with infinite elements to its reliable neighbor $i$. Another example is that Byzantine agent $b$ can deter all reliable neighbors from achieving consensus at iteration $k$, via sending $\tilde{x}_{ib,k}$ with various values to its different reliable neighbors $i\in{\mathcal{R}_{b}}$. The primary cause of the aforementioned issues lies in the vulnerability of the aggregation step (5b) to Byzantine agents. In fact, similar security threats also prevail in decentralized work [12, 10, 1, 13, 14, 15, 11, 2]. To address this issue, the SGD family receives two important extensions, RSA [28] and [27], both of which achieve Byzantine resilience based on a resilient aggregation mechanism [36], where [27] is a decentralized extension of RSA [28]. The theoretical analysis of both RSA [28] and [27] is based on a bounded-variance assumption on the local stochastic gradient. With this assumption and the other standard assumptions (see [27] for details), the sequence ${\left\{{{x_{k}}}\right\}_{k\geq 0}}$ generated by the decentralized algorithm proposed in [27] takes a convergent form of
@@
where $\eta$ is a positive constant satisfying $10<\eta{\alpha_{k}}<$ 1, ${\Delta_{0}}:=\sum\nolimits_{i\in\mathcal{R}}{32n{\phi^{2}}{{\left|{{\mathcal{R}_{i}}}\right|}^{2}}}+4n{\phi^{2}}{\left|{{\mathcal{B}_{i}}}\right|^{2}}+2\sigma_{i}^{2}$ (${\sigma_{i}}>0$ is the bounded variance yielded by the biased estimation of the local batch gradients) and ${\Delta_{1}}:=\left({n{\phi^{2}}/\gamma}\right)\sum\nolimits_{i\in{\cal R}}{{{\left|{{{\cal B}_{i}}}\right|}^{2}}}$. Based on (6), one can establish either a sub-linear convergence rate with a smaller convergence error determined by the number of Byzantine agents, or a faster linear convergence rate with a larger convergence error determined jointly by the number of Byzantine agents and the bounded variance. In fact, this bounded variance ($\sigma_{i}^{2}$) exists commonly in recent literature, such as [32, 33, 27, 14, 31, 26, 24]. Therefore, this paper aims to asymptotically reduce this bounded variance in the linear convergence case and circumvent the bounded-variance assumption. Inspired by the recent exploration of decentralized VR stochastic gradient algorithms diffusion-AVRG [2], S-DIGing [15], GT-SAGA/GT-SVRG [11], and GT-SARAH [1] that seek the solution to an optimization problem under an ideal Byzantine-free situation, we introduce two popular localized variance-reduced techniques SAGA [43] and LSVRG [44] to asymptotically reduce the noise variance in local gradient estimation. These two VR techniques allow us to derive a unified theoretical result on Byzantine-resilient decentralized stochastic proximal-gradient optimization.
@@ III-B A General Algorithmic Framework
Based on the above analysis, we propose a Byzantine-resilient decentralized stochastic-gradient algorithmic framework in Algorithm 1 to resolve (4) in the presence of Byzantine agents. Note that we denote temporarily the local stochastic gradient by $r_{i,k}$, which will be specified in Steps 4-5 of Algorithms 2-3.
Input: a proper constant or decaying step-size $\alpha_{k}>0$.
Initialize: an arbitrary starting point $x_{i,0}\in{\mathbb{R}^{n}}$ and a proper penalty parameter ${\phi}>0$.
1 for $k=0,1,\ldots,$ do
2       for each reliable agent $i\in\mathcal{R}$ do
3             Transmit its current local model ${x_{i,k}}$ to its neighbors $j\in{\mathcal{N}_{i}}$ and receive the true information ${x_{j,k}}$ or untrue information ${z_{ij,k}}$ from its neighbors;
4             Evaluate the local stochastic gradient $r_{i,k}$;
5             Calculate an intermediate variable according to the local resilient stochastic subgradient descent step:
${\bar{x}_{i,k}}={x_{i,k}}-{\alpha_{k}}r_{i,k}-{\alpha_{k}}{\phi}\sum\limits_{j\in{\mathcal{N}_{i}}}{{\partial_{{x_{i}}}}{{\left\|{{x_{i,k}}-v_{ij,k}}\right\|}_{a}}},$
where ${v_{ij,k}}:=\left\{\begin{gathered}{x_{j,k}},{\text{if}}\;j\in{\mathcal{R}_{i}}\hfill\\ {z_{ij,k}},{\text{if}}\;j\in{\mathcal{B}_{i}}\hfill\\ \end{gathered}\right..$
6             Update its current local model according to the local proximal gradient step:
${x_{i,k+1}}=\arg\mathop{\min}\limits_{\tilde{x}\in{\mathbb{R}^{n}}}\left\{{g\left({\tilde{x}}\right)+\frac{1}{{2{\alpha_{k}}}}\left\|{\tilde{x}-{{\bar{x}}_{i,k}}}\right\|_{2}^{2}}\right\}.$
7      for each Byzantine agent $i\in\mathcal{B}$ do
8            Send an arbitrary vector ${z_{ij,k}}$ to its neighbor $j$, $j\in{\mathcal{N}_{i}}$.
9
Output: all decision variables ${x_{i,k}}$, $i\in\mathcal{V}$, until a prescribed criterion is satisfied.
Algorithm 1 Prox-DBRO-VR Framework.